Determine the amount of each of the two solutions that must be mixed to make 250 mL of a 40% acid solution. Round to the nearest milliliter.
1.) 25% acid, 80% acid
2.) 30% acid, 60% acid
3.) 15% acid, 75% acid
4.) 10% acid, 45% acid

Answer by TimothyLamb(4379)

I will do the first, you do the other three using the same method:
---
1) 25% acid, 80% acid
---
x = mL of first solution
y = mL of second solution
---
x + y = 250
0.25x + 0.80y = 0.40*250
---
put the system of linear equations into standard form
---
x + y = 250
0.25x + 0.80y = 100
---
copy and paste the above standard form linear equations in to this solver:
http://sooeet.com/math/system-of-linear-equations-solver.php
---
answer:
x = 181.81818
y = 68.181818
---
x = 182
y = 68
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---